# Decision-Making Behavior Evaluation Framework for LLMs under Uncertain Context

### v1

@@ 1 Introduction
In recent years, the deployment of large language models (LLMs) such as ChatGPT-4.0-Turbo [25], Claude-3-Opus [26], and Gemini-1.0-pro [27] has revolutionized various fields by providing sophisticated, human-like responses to a multitude of queries. Their applications span from answering everyday questions and content generation to complex decision-support systems in healthcare, finance, and beyond [28, 29, 30]. Domain-specific LLMs have also emerged, serving as customer service agents, investment assistants, and more [31, 32, 33]. Understanding their internal decision-making tendencies becomes crucial as these models become increasingly integral to decision-making processes. How do LLMs handle risk and uncertainty in comparison to human decision-makers? To what extent do LLMs exhibit biases when socio-demographic features are introduced into their decision-making processes [34]? Can we trust LLMs to make fair and ethical decisions across diverse contexts and populations? All of these questions deserve careful investigation and confident answers.
@@
Given the prominence of LLMs in aiding decision-making, it is imperative to assess whether their behaviors align with or diverge from human tendencies. Previous empirical studies have examined the rationality and social behaviors of LLMs, yet their intrinsic decision-making processes remain insufficiently understood [14, 16], especially when the LLMs are given contextual prompts about the identities or characteristics of the user.
In this study, we experiment with three feature assignments: a baseline with no demographic context (context-free), human-like demographic distributions, and augmented distributions with minority group characteristics. We assess how LLMs’ decision-making process aligns with or diverges from human-like behavior and uncovers potential biases and ethical concerns, emphasizing the consideration for fairness in deploying these models across diverse user groups. To summarize, our contributions are threefold:
@@
    We apply this framework to three state-of-the-art commercial LLM models: ChatGPT-4.0-Turbo, Claude-3-Opus, and Gemini-1.0-pro, to assess their decision-making behavior in a context-free setting. Our results indicate that LLMs generally exhibit human-like patterns: risk aversion, loss aversion, and overweighting small probabilities. Nevertheless, there are significant variations in the degree to which these behaviors are expressed across different LLM.
@@ 2.1 LLMs’ Behavior Study
Numerous studies within social science have evaluated the alignment between LLM and human behaviors and decision-making processes.
For instance, research has shown that LLMs demonstrate economic behaviors, including adherence to downward-sloping demand curves, diminishing marginal utility, status quo bias, and the endowment effect [10, 9, 12]. LLMs also show consistency with human-like behavior patterns in the psychology binary moral judgment experiment [11].
@@
Another critical issue is the fairness and bias in LLM processing, particularly concerning specific personality traits. Work[16] has shown that ChatGPT-3.5 and ChatGPT-4 exhibit significant biases and a decline in performance when handling information pertaining to minority groups. This underscores the need to develop a framework that overcomes existing limitations and quantitatively evaluates LLM behavior under uncertainty, exploring the effect and bias related to human demographic features.
@@ 3 Preliminary
$w(p)$ $=\exp[-(\ln p)^{\alpha}]$
where $x$, $y$ are experiment outcomes, and $p$, $q$ are the associated probabilities. Probabilities are weighted by $w(p)$ and $w(q)$. $\sigma$ captures the curvature of the value function in which the person is risk-seeking if $\sigma<0$, risk-neutral if $\sigma=0$, and risk-averse if $\sigma>0.$ High $\lambda$ implies a more loss averse person; $\alpha$ determines the weighting of choices of different risk and outcome. $\alpha<1$ indicates an overweighting of big gains or losses. When $\lambda=\alpha=1$, the model reduces to the expected utility theory. The interpretation of the three parameters $\sigma$, $\alpha$, and $\lambda$ are illustrated in Figure 1.
@@ 4 Framework and Design
Step 1: Experimentation Design Three series of multiple-choice-list experiments, as shown in Appendix B, are used to elicit the decision-making preferences of the subjects. Each experiment involves presenting subjects with a series of choices between different probabilistic outcomes, also known as lottery games. In each lottery game, subjects are asked to make choices between the options with varying probabilities and outcomes, allowing researchers to infer their preferences and behavioral patterns. Each series is designed to test different aspects of decision-making under uncertainty: Series 1 and 2 focus on positive outcomes to determine the effects of $\sigma$ and $\alpha$; and Series 3 introduces negative outcomes to evaluate the parameter $\lambda$ for loss aversion.
@@
$v(x)=x^{1-\sigma},\quad w(p)=\exp[-(\ln p)^{\alpha}].$
@@
$v(y_{A_{n}})+w(p_{A_{n}})(v(x_{A_{n}})-v(y_{A_{n}}))$ $=y_{A_{n}}^{1-\sigma}+\exp[-(\ln p_{A_{n}})^{\alpha}](x_{A_{n}}^{1-\sigma}-y_{A_{n}}^{1-\sigma}),$
$v(y_{B_{n}})+w(p_{B_{n}})(v(x_{B_{n}})-v(y_{B_{n}}))$ $=y_{B_{n}}^{1-\sigma}+\exp[-(\ln p_{B_{n}})^{\alpha}](x_{B_{n}}^{1-\sigma}-y_{B_{n}}^{1-\sigma}),$
$\Rightarrow\quad y_{A_{n}}^{1-\sigma}+\exp[-(\ln p_{A_{n}})^{\alpha}](x_{A_{n}}^{1-\sigma}-y_{A_{n}}^{1-\sigma})$ $>y_{B_{n}}^{1-\sigma}+\exp[-(\ln p_{B_{n}})^{\alpha}](x_{B_{n}}^{1-\sigma}-y_{B_{n}}^{1-\sigma}),$
@@
$v(y_{A_{n+1}})+w(p_{A_{n+1}})(v(x_{A_{n+1}})-v(y_{A_{n+1}}))$ $<v(y_{B_{n+1}})+w(p_{B_{n+1}})(v(x_{B_{n+1}})-v(y_{B_{n+1}})),$
$\Rightarrow\quad y_{A_{n+1}}^{1-\sigma}+\exp[-(\ln p_{A_{n+1}})^{\alpha}](x_{A_{n+1}}^{1-\sigma}-y_{A_{n+1}}^{1-\sigma})$ $<y_{B_{n+1}}^{1-\sigma}+\exp[-(\ln p_{B_{n+1}})^{\alpha}](x_{B_{n+1}}^{1-\sigma}-y_{B_{n+1}}^{1-\sigma}).$
@@ 5.1.1 Results and Key Findings
We rank the models according to three dimensions of decision-making behavior under uncertainty, as shown in Figure 3, raw data is available in the Appendix Table 5. Interestingly, all three models share a unanimous preference for risk-averse decision-making tendencies. Across these LLMs, the average $\sigma$ value exceeds 0, and even their minimum $\sigma$ value remains greater than 0. The TCN model allows LLMs to exhibit diverse risk preferences, but despite this flexibility, our findings still reveal a unanimous preference for risk-averse behavior among LLMs, which is consistent with human behavior to a certain extent. ChatGPT leans towards conservative choices with higher rewards, as reflected in its higher risk aversion ($\sigma$) but it shows the lowest concern for potential losses ($\lambda$). With probability weighting parameter $\alpha>1$, it diverges from human norms by favoring unlikely events or outcomes. Conversely, Claude adopts a riskier approach, with lower risk aversion, but has higher loss aversion, and probability underweighting. Gemini balances risk and caution, exhibiting moderate risk-taking tendencies, loss aversion, and balanced probability weighting behavior. An uncommon less-than-1 $\alpha$ of ChatGPT could significantly influence decision-making processes in various applications: (1) Dialogue Systems: It might produce more creative or unexpected responses, potentially enhancing user engagement by introducing novel ideas or perspectives. (2) Content Generation: The model may favor generating content that explores less common scenarios or narratives, offering a fresh take but with an increased risk of deviating from expected norms.
@@ 5.2 Embedded Demographic Features
below lower secondary, lower secondary, upper secondary, short-cycle tertiary, bachelor, and graduate degrees
@@
Table 2: The 38 Personas across 10 socio-demographic groups that we explore in this ttudy.
In addition to the context-free decision-making baseline, we investigate potential variations in decision patterns among LLMs based on embedded demographic features, including gender, age, education level, marital status, and living area. Given the existing literature highlighting biases within LLMs when encountering such features [16], we introduce an augmentation technique that involves randomly incorporating other minority features into the basic demographic framework. These features encompass sexual orientation, disability, religion, race, and political affiliation. The augmentation of our demographic profiles results in 10 distinct socio-demographic groups as outlined in Table 2. We first generate a random distribution of demographic profiles by giving LLMs foundational demographic features randomly and recording their answers. We then apply the real-world distribution across the countries and document the results, where the statistical data are from the World Bank dataset [37].
@@ 5.2.1 Results and Key Findings
ChatGPT shifts towards more riskier decision-making with lower risk aversion while maintaining consistent loss aversion. Notably, $\alpha$ shifts from small probability underweighting to overweighting, aligning more closely with typical human behavior. Claude shows stable risk preferences and probability weighting but reduced loss aversion, aligning more closely with human tendencies. Gemini becomes more conservative in risk-taking and shows heightened loss aversion with demographic embedding. Contrary to ChatGPT, this model shifts from $\alpha$ smaller than 1 to small probability underweighting.
@@
a. Foundational Demographic Information: We use Ordinary Least Squares (OLS) regression to explore the determinants of decision-making behaviors under uncertainty, using $\sigma$, $\alpha$, and $\lambda$ as dependent variables and foundational demographic features as independent variables. We encode categorical features with binaries. We then have the regression model below, where $\beta_{n}$ are coefficients for demographic variables and $\beta_{0}$ is the intercept. $n$ denotes $n$-th feature, $i$ denotes the $i$-th observation, and $\epsilon$ denotes the error offset in regression.
$\text{RiskParameter}_{i}=\beta_{0}+\beta_{n}\text{Demographic}_{ni}+\epsilon_{i}$
@@
Younger Individuals (<25 years): Claude shows significant higher tendency to overweight
@@
Claude may think married people are more likely to overweight small probabilities.
@@
Table 3 highlights key findings from this regression study. As an example of features that have a significant impact on these parameters, we graph comparisons of the average parameter values in Age Impact and Education Level in Figure 5(a) and 5(b). Complete regression results are presented in Figure 6 and Table 7 in Appendix A. Incorporated with demographic features, the average parameters across all three models show distinct differences. Additionally, each model exhibits unique sensitivity to the features that affect its application in various demographic contexts.
@@
Claude demonstrates a broader sensitivity to demographic variables in both loss aversion and probability weighting, particularly for young and rural populations. It is also sensitive to education level and marital status. This model’s extensive demographic responsiveness could enhance its adaptability in diverse settings but also introduce the risk of unfairness. For ChatGPT and Gemini, the analysis indicates a generally lower sensitivity to demographic variables, consistent across diverse user groups. ChatGPT exhibits a significant gender difference in risk preferences, where females show reduced risk aversion. Gemini shows strong effects on loss aversion in the younger age group and lower education level group.
@@
$\text{RiskParameter}_{i}=\beta_{0}+\beta_{n}\text{Demographic}_{ni}+\gamma_{m}\text{Advanced Demographic}_{mi}+\epsilon_{i}$
Table 4 highlights key findings from this regression study. Similarly, as an example, we graph comparisons of the average parameter values in Religious Background and Sexual Orientation in Figure 5(c) and 5(d). Complete regression results are presented in Figure 7 and Table 8 in Appendix A. The distinct responses patterns from advanced features embedded LLMs may suggest unique capabilities or potential biases. Claude demonstrates broader adaptability to various demographic factors, including sexual orientation, ethnicity, disability, religious background, and political beliefs. ChatGPT exhibits a pronounced sensitivity to political beliefs in terms of loss aversion. This model also shows targeted sensitivity toward physically disabled individuals in probability weighting. Gemini, meanwhile, displays a specific sensitivity to ethnicity, particularly in loss aversion for African and Hispanic groups. Unlike the other models, it maintains a politically neutral stance when making decisions.

### v2

@@ 1 Introduction
In recent years, the deployment of large language models (LLMs) such as ChatGPT-4.0-Turbo [37], Claude-3-Opus [12], and Gemini-1.0-pro [33] has revolutionized various fields by providing sophisticated, human-like responses to a multitude of queries [6, 41, 25, 40, 11, 26]. Their applications span from answering everyday questions and content generation to complex decision-support systems in healthcare, finance, and beyond [31, 8, 38]. Domain-specific LLMs have also emerged, serving as customer service agents, investment assistants, and more [29, 23, 17]. Understanding their internal decision-making tendencies becomes crucial as these models become increasingly integral to decision-making processes. How do LLMs handle risk and uncertainty in comparison to human decision-makers? To what extent do LLMs exhibit biases when socio-demographic features are introduced into their decision-making processes [20]? Can we trust LLMs to make fair and ethical decisions across diverse contexts and populations? All of these questions deserve careful investigation and confident answers.
@@
Given the prominence of LLMs in aiding decision-making, it is imperative to assess whether their behaviors align with or diverge from human tendencies. Previous empirical studies have examined the rationality and social behaviors of LLMs, yet their intrinsic decision-making processes remain insufficiently understood [13, 15], especially when the LLMs are given contextual prompts about the identities or characteristics of the user. In this study, we experiment with three feature assignments: a baseline with no demographic context (context-free), human-like demographic distributions, and augmented distributions with minority group characteristics. We assess how LLMs’ decision-making process aligns with or diverges from human-like behavior and uncovers potential biases and ethical concerns, emphasizing the consideration for fairness in deploying these models across diverse user groups. To summarize, our contributions are threefold:
@@
    We apply this framework to three state-of-the-art commercial LLM models: ChatGPT-4.0-Turbo, Claude-3-Opus, and Gemini-1.0-pro, to assess their decision-making behavior in a context-free setting. Our results indicate that LLMs generally exhibit human-like patterns: risk aversion, loss aversion, and overweighting small probabilities. Nevertheless, there are significant variations in the degree to which these behaviors are expressed across different LLMs.
@@ 2.1 LLMs’ Behavior Study
Numerous studies within social science have evaluated the alignment between LLM and human behaviors and decision-making processes. For instance, research has shown that LLMs demonstrate economic behaviors, including adherence to downward-sloping demand curves, diminishing marginal utility, status quo bias, and the endowment effect [5, 19, 7]. LLMs also show consistency with human-like behavior patterns in the psychology binary moral judgment experiment [9].
@@
Another critical issue is the fairness and bias in LLM processing, particularly concerning specific demographic or personality traits. Work[15] has shown that ChatGPT-3.5 and ChatGPT-4 exhibit significant biases and a decline in performance when handling information pertaining to minority groups in humans. This underscores the need to develop a framework that overcomes existing limitations and quantitatively evaluates LLM behavior under uncertainty, exploring the effect and bias related to human demographic features.
@@ 3 Preliminary
$w(p)$ $=\exp[-(-\ln p)^{\alpha}]$ (3)
where $x$, $y$ are experiment outcomes, and $p$, $q$ are the associated probabilities. Probabilities are weighted by $w(p)$ and $w(q)$. $\sigma$ captures the curvature of the value function in which the person is risk-seeking if $\sigma<0$, risk-neutral if $\sigma=0$, and risk-averse if $\sigma>0.$ High $\lambda$ implies a more loss averse person; $\alpha$ determines the weighting of choices of different risk and outcome. $\alpha<1$ indicates an overweighting of small probabilities. When $\lambda=\alpha=1$, the model reduces to the expected utility theory. The interpretation of the three parameters $\sigma$, $\alpha$, and $\lambda$ are illustrated in Figure 1.
@@ 4 Framework and Design
Step 1: Experimentation Design Three series of multiple-choice-list experiments, as shown in Appendix B, are used to elicit the decision-making preferences of the subjects. Each experiment involves presenting subjects with a series of choices between different probabilistic outcomes, also known as lottery games. In each lottery game, subjects are asked to make choices between the options with varying probabilities and outcomes, allowing researchers to infer their preferences and behavioral patterns. Each series is designed to test different aspects of decision-making under uncertainty: Series 1 and 2 focus on positive outcomes to determine the effects of $\sigma$ and $\alpha$; and Series 3 introduces negative outcomes to evaluate the parameter $\lambda$ for loss aversion. In our experiment design, we treat LLMs similarly to human populations, where each query to the LLM represents an individual interaction, analogous to testing a different human subject. Just as individual humans have fixed risk preferences in a given context, we assume fixed parameters for each interaction with the LLM. By repeating these interactions across multiple queries, we capture the overall behavioral tendencies of the LLM, just as repeated studies in human populations help reveal broader trends.
@@
$v(x)=x^{1-\sigma},\quad w(p)=\exp[-(-\ln p)^{\alpha}].$
@@
$v(y_{A_{n}})+w(p_{A_{n}})(v(x_{A_{n}})-v(y_{A_{n}}))$ $=y_{A_{n}}^{1-\sigma}+\exp[-(-\ln p_{A_{n}})^{\alpha}](x_{A_{n}}^{1-\sigma}-y_{A_{n}}^{1-\sigma}),$ (4)
$v(y_{B_{n}})+w(p_{B_{n}})(v(x_{B_{n}})-v(y_{B_{n}}))$ $=y_{B_{n}}^{1-\sigma}+\exp[-(-\ln p_{B_{n}})^{\alpha}](x_{B_{n}}^{1-\sigma}-y_{B_{n}}^{1-\sigma}),$ (5)
$\Rightarrow\quad y_{A_{n}}^{1-\sigma}+\exp[-(-\ln p_{A_{n}})^{\alpha}](x_{A_{n}}^{1-\sigma}-y_{A_{n}}^{1-\sigma})$ $>y_{B_{n}}^{1-\sigma}+\exp[-(-\ln p_{B_{n}})^{\alpha}](x_{B_{n}}^{1-\sigma}-y_{B_{n}}^{1-\sigma}),$ (6)
@@
$v(y_{A_{n+1}})+w(p_{A_{n+1}})(v(x_{A_{n+1}})-v(y_{A_{n+1}}))$ $<v(y_{B_{n+1}})+w(p_{B_{n+1}})(v(x_{B_{n+1}})-v(y_{B_{n+1}})),$ (7)
$\Rightarrow\quad y_{A_{n+1}}^{1-\sigma}+\exp[-(-\ln p_{A_{n+1}})^{\alpha}](x_{A_{n+1}}^{1-\sigma}-y_{A_{n+1}}^{1-\sigma})$ $<y_{B_{n+1}}^{1-\sigma}+\exp[-(-\ln p_{B_{n+1}})^{\alpha}](x_{B_{n+1}}^{1-\sigma}-y_{B_{n+1}}^{1-\sigma}).$ (8)
@@ 5.1.1 Results and Key Findings
We rank the models according to three dimensions of decision-making behavior under uncertainty, as shown in Figure 3, raw data is available in the Appendix Table 5. Interestingly, all three models share a unanimous preference for risk-averse decision-making tendencies. Across these LLMs, the average $\sigma$ value exceeds 0, and even their minimum $\sigma$ value remains greater than 0. The TCN model allows LLMs to exhibit diverse risk preferences, but despite this flexibility, our findings still reveal a unanimous preference for risk-averse behavior among LLMs, which is consistent with human behavior to a certain extent. ChatGPT leans towards conservative choices with higher rewards, as reflected in its higher risk aversion ($\sigma$) but it shows the lowest concern for potential losses ($\lambda$). With probability weighting parameter $\alpha>1$, it diverges from human norms. Conversely, Claude adopts a riskier approach, with lower risk aversion, but has higher loss aversion, and small-probability overweighting. Gemini balances risk and caution, exhibiting moderate risk-taking tendencies, loss aversion, and balanced probability weighting behavior. An uncommon greater-than-1 $\alpha$ suggests that ChatGPT would perceive unlikely events as even less likely than they are. This could have the following implications: (1) Dialogue Systems: It may produce more conservative responses, which might make it better suited for providing safe, predictable information.(2) Content Generation: It may avoid rare scenarios, leading to content that aligns more with conventional or high-probability outcomes, which could reduce risk but also the potential for novelty and creativity.
@@ 5.2 Embedded Demographic Features
below lower secondary, lower secondary, upper secondary,
short-cycle tertiary, bachelor, and graduate degrees
@@
Table 2: The Personas across 10 socio-demographic groups that we explore in this study.
In addition to the context-free decision-making baseline, we investigate potential variations in decision patterns among LLMs based on embedded demographic features, including gender, age, education level, marital status, and living area. Given the existing literature highlighting biases within LLMs when encountering such features [15], we introduce an augmentation technique that involves randomly incorporating other minority features into the basic demographic framework. These features encompass sexual orientation, disability, religion, race, and political affiliation. The augmentation of our demographic profiles results in 10 distinct socio-demographic groups as outlined in Table 2. These personas provide a comprehensive representation of the diverse demographic landscape under examination. We first generate a distribution of demographic profiles by randomly assigning LLMs foundational demographic features and recording their answers. We then apply the real-world distribution across the countries to reflect realistic demographics of human communities and document the results, where the statistical data are from the World Bank dataset [36].
@@ 5.2.1 Results and Key Findings
ChatGPT shifts towards much riskier decision-making with lower risk aversion ($\sigma$), while maintaining consistent loss aversion ($\lambda$). Notably, $\alpha$ shifts from small probability underweighting to overweighting, aligning more closely with typical human behavior. Claude shows stable risk preferences and probability weighting but reduced loss aversion, aligning more closely with human tendencies. Gemini becomes more conservative in risk-taking and shows heightened loss aversion with demographic embedding. Contrary to ChatGPT, this model shifts from small probability overweighting to underweighting with $\alpha$ slightly higher than 1.
@@
a. Foundational Demographic Information: We use Ordinary Least Squares (OLS) regression to explore the determinants of decision-making behaviors under uncertainty, using $\sigma$, $\alpha$, and $\lambda$ as dependent variables and foundational demographic features as independent variables. We encode categorical features with binaries. For example, "Age < 25 years" is set to 1 for individuals younger than 25 years old and 0 otherwise. We then have the regression model below, where $\beta_{n}$ are coefficients for demographic variables and $\beta_{0}$ is the intercept. $n$ denotes $n$-th feature, $i$ denotes the $i$-th observation, and $\epsilon$ denotes the error offset in regression.
$\text{RiskParameter}_{i}=\beta_{0}+\beta_{n}\text{Demographic}_{ni}+\epsilon_{i}$ (9)
@@
Younger Individuals (<25 years): Claude shows a significantly higher tendency to overweight
@@
Claude may think married people are less likely to be risk-averse.
@@
Table 3 highlights key findings from this regression study. As an example of features that have a significant impact on these parameters, we graph comparisons of the average parameter values in Age Impact and Education Level in Figure 5(a) and 5(b). The coefficients are extracted from the complete regression as presented in Figure 6. The raw data are also included in Table 7 in Appendix A. Incorporated with demographic features, the average parameters across all three models show distinct differences. Additionally, each model exhibits unique sensitivity to the features that affect its application in various demographic contexts.
Figure 6: Influence of Fundamental Demographic Feature: Estimated Coefficients
In the above examples, Claude demonstrates a broader sensitivity to demographic variables in both loss aversion and probability weighting, particularly for young and rural populations. It is also sensitive to education level and marital status. This model’s extensive demographic responsiveness could enhance its adaptability in diverse settings but also introduce the risk of unfairness. For ChatGPT and Gemini, the analysis indicates a generally lower sensitivity to demographic variables, consistent across diverse user groups. ChatGPT exhibits a significant gender difference in risk preferences, in which females show reduced risk aversion than males. Gemini shows strong effects on loss aversion in the younger age group and lower education level group.
@@
$\text{RiskParameter}_{i}=\beta_{0}+\beta_{n}\text{Demographic}_{ni}+\gamma_{m}\text{Advanced Demographic}_{mi}+\epsilon_{i}$ (10)
In examining the additional advanced demographic sensitivities of LLMs, distinct patterns of behavior highlight their unique capabilities and potential biases. Table 4 highlights key findings from this regression study. Similarly, as an example, we graph comparisons of the average parameter values in Religious Background and Sexual Orientation in Figure 5(c) and 5(d). Complete regression results and raw data are presented in Figure 7 and Table 8 in Appendix A. The distinct response patterns from advanced features-embedded LLMs may suggest unique capabilities or potential biases. Claude demonstrates broader adaptability to various demographic factors, including sexual orientation, ethnicity, disability, religious background, and political beliefs.
In these two examples, ChatGPT exhibits a pronounced sensitivity to political beliefs in terms of loss aversion. This model also shows targeted sensitivity toward physically disabled individuals in probability weighting, being more likely to overweight small probability events. Meanwhile, Gemini displays a specific sensitivity to ethnicity, particularly in loss aversion for African and Hispanic groups. Unlike the other models, it maintains a politically neutral stance when making decisions.